(a) Show that if yp(t) is a solution of duydu 2 + +qy = g(t), then kyp(t) is a solution ofdy the +Pdt peop + qv = kg (0 -+qy = kg(t) an2x for any constant k. (b) Find the general solution of d²v - + 8y = 5 cost. d12
Answer #1

Sol? @ Since Ypct) is a solution of deporte + palete + ay = g(t) , therefore e n + papie + 2 dp = get) 0 Let Kyplt) = m(t), tI rzaysa (778035 = R(8 +09+0) 78975 = Rp + Hoy te te gute Hence, ky plt) is solution of het hele pode 2 + 2y = kg (t). (7.78*- Auxilary equation is mat6m +8 =0 (m+4) (m+2) =0 m = -2,-4 Complementary solution (say ye) is Tyc = qe-2t + ca exht I whereSo, general solution (sayy) of given differential equation is y = Yc + Yp y = cie-2t +gecut + (76o8t + 6 sint) where calca ar